OpenStudy (anonymous):

6. which equation would be used to generate the data in the following table ? x / y 1 / 8 2 / 16 3 / 24 4 / 32 ( 1point ) A. x = 2y b. y = x + 7 c. y = 2x d. y = 8x

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

x = 2y Take a point, let's use (1, 8) 1 = 2(8) 1 = 16 Incorrect. y = x + 7 8 = 1 + 7 8 = 8 Correct. Let's try another point: (2, 16) y = x + 7 16 = 2 + 7 16 = 9 Incorrect. y = 2x (1, 8) 1 = 2(8) 1 = 16 Incorrect. y = 8x (1, 8) 8 = 8(1) 8 = 8 Correct. So your answer is D.